# thanks: Equal contributionthanks: Equal contribution

### v1

@@ Appendix C Generalized consumer-resource models
$R_{\alpha/pm}(t)\simeq R_{\alpha/pm}(0)\,,\quad N_{p}(t)+N_{m}(t)\simeq N_{0}\,.$ (105)
@@
$\frac{df}{dt}=\,$ $f(1-f)\left(g_{m}\left(\vec{R}_{/pm}(0)\right)-g_{p}\left(\vec{R}_{/pm}(0)\right)-m_{m}+m_{p}\right$. (106)
@@
$=$ $\,f(1-f)\left(s-\eta(f-f_{0})\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (108)
where the invasion fitness difference is
$s=$ $\,g_{m}\left(\vec{R}_{/pm}(0)\right)-g_{p}\left(\vec{R}_{/pm}(0)\right)-m_{m}+m_{p}$ (109)
$\,-N_{0}\sum_{\alpha\beta}(C_{m\alpha}-C_{p\alpha})\left(\left(I-P\right)Q^{-1}\right)_{\alpha\beta}\left(E_{p\beta}(1-f_{0})+E_{m\beta}f_{0}\right)$ (110)
@@ Appendix D Derivation and analysis of fixation probability
$\frac{df}{dt}=f(1-f)\left(s-\eta(f-f_{0})\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (112)
where $f_{0}$ is the initial frequency, $s$ is the initial difference in invasion fitness, $\eta$ is the strength of ecological effects depending on details of the ecological model, and $N_{\mathrm{eff}}$ is the initial effective population size of the parent and the mutant.
@@
$\frac{\partial p}{\partial t}=x(1-x)(s-\eta(x-f_{0}))\frac{\partial p}{\partial x}+\frac{x(1-x)}{2N_{\mathrm{eff}}}\frac{\partial^{2}p}{\partial x^{2}}\,.$ (113)
@@
$0=x(1-x)(s-\eta(x-f_{0}))\frac{\partial p^{*}}{\partial x}+\frac{x(1-x)}{2N_{\mathrm{eff}}}\frac{\partial^{2}p^{*}}{\partial x^{2}}\,$, (114)
$\Rightarrow\frac{\partial p^{*}}{\partial x}\propto\exp\left[N_{\mathrm{eff}}\left(\eta x^{2}-2(s+\eta f_{0})x\right)\right]\,.$ (115)
@@
$=\frac{\mathrm{Erfi}\left(\alpha(\tilde{s}+f_{0})\right)-\mathrm{Erfi}\left(\alpha\tilde{s}\right)}{\mathrm{Erfi}\left(\alpha(\tilde{s}+f_{0})\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}-1+f_{0})\right)}\,$, (117)
@@
is the imaginary error function. We have denoted $\alpha=\sqrt{N_{\mathrm{eff}}\eta}$ as the ratio of strength between ecological effects and demographic noise, and $\tilde{s}=s/\eta$ is proportional to the “dressed invasion fitness” [17].
@@
For example, when $s=0$,
$p_{\mathrm{fix}}(s=0)=\frac{\mathrm{Erfi}\left(\alpha f_{0}\right)}{\mathrm{Erfi}\left(\alpha f_{0}\right)+\mathrm{Erfi}\left(\alpha(1-f_{0})\right)}\,.$ (120)
@@
$p_{\mathrm{fix}}(s=0,\alpha\ll 1)\simeq\frac{\alpha f_{0}}{\alpha f_{0}+\alpha(1-f_{0})}=f_{0}\,.$ (121)
@@
$p_{\mathrm{fix}}(s=0,\alpha\gg 1)\sim e^{-\alpha^{2}}\,.$ (122)
@@
We see that $p_{\mathrm{fix}}(s=0)$ matches Kimura’s prediction when $\alpha\ll 1$, but exponentially suppressed by $\alpha$ otherwise.
@@
The fixation probability for neutral mutations suggests that there is a crossover between the regimes of weak and strong ecological effects, characterized by small and large $\eta$ respectively. First, for small $\eta$ with $\alpha\ll 1$, we have $|\alpha\tilde{s}|\gg 1$ for any finite $s$ since $\alpha\tilde{s}$ scales as $\eta^{-1/2}$. Therefore, we can use the exponential approximation for all the terms and $p_{\mathrm{fix}}$ becomes
$p_{\mathrm{fix}}$ $\simeq\frac{\exp\left(\alpha^{2}(\tilde{s}+f_{0})^{2}\right)-\exp\left(\alpha^{2}\tilde{s}^{2}\right)}{\exp\left(\alpha^{2}(\tilde{s}+f_{0})^{2}\right)-\exp\left(\alpha^{2}(\tilde{s}-1+f_{0})^{2}\right)}$ (123)
$\simeq\frac{1-\exp(-2N_{\mathrm{eff}}sf_{0})}{1-\exp(-2N_{\mathrm{eff}}s)}\,$, (124)
@@
which is exactly Kimura’s formula for fixation probability. Combining the result for neutral mutation, we see that Kimura’s formula works equally well in ecological context as long as $\alpha\ll 1$. On the other hand, for large $\eta$ with $\alpha\gg 1$, the behavior of $p_{\mathrm{fix}}$ is divided into different regimes according to $s$. Now for $|\alpha\tilde{s}|\ll 1$, we again have
$p_{\mathrm{fix}}\simeq\frac{\mathrm{Erfi}\left(\alpha f_{0}\right)}{\mathrm{Erfi}\left(\alpha f_{0}\right)+\mathrm{Erfi}\left(\alpha\right)}\sim e^{-\alpha^{2}}\,.$ (125)
For $|\alpha\tilde{s}|\gg 1$, we have
$p_{\mathrm{fix}}$ $\simeq\frac{\exp\left(\alpha^{2}(\tilde{s}+f_{0})^{2}\right)-\exp\left(\alpha^{2}\tilde{s}^{2}\right)}{\exp\left(\alpha^{2}(\tilde{s}+f_{0})^{2}\right)-\mathrm{sgn}(\tilde{s}-1+f_{0})\exp\left(\alpha^{2}(\tilde{s}-1+f_{0})^{2}\right)}$ (126)
$=\frac{1-\exp\left(-\alpha^{2}f_{0}(2\tilde{s}+f_{0})\right)}{1-\mathrm{sgn}(\tilde{s}-1+f_{0})\exp\left(-\alpha^{2}(2\tilde{s}-1+2f_{0})\right)}\,.$ (127)
@@
The behavior of the above expression is further divided into two regimes. When $\tilde{s}\ll 1/2-f_{0}$, the denominator is dominating and $p_{\mathrm{fix}}$ is still exponentially suppressed:
$p_{\mathrm{fix}}\sim\exp\left(-2N_{\mathrm{eff}}(s-\eta(1/2-f_{0}))\right)\,.$ (128)
@@
When $\tilde{s}\gg 1/2-f_{0}$, the denominator becomes very close to $1$ and the numerator becomes dominating instead. We then have
$p_{\mathrm{fix}}\simeq\frac{1-\exp(-2N_{\mathrm{eff}}sf_{0})}{1-\mathrm{sgn}(\tilde{s}-1+f_{0})\exp(-2N_{\mathrm{eff}}s)}\,$, (129)
@@
which matches Kimura’s formula approximately for $\tilde{s}<1-f_{0}$ and exactly for $\tilde{s}>1-f_{0}$.
@@
In conclusion, while $p_{\mathrm{fix}}$ for negative $s$ is always exponentially suppressed regardless of ecology, the presence of a large $\eta$ now also exponentially suppresses $p_{\mathrm{fix}}$ with positive $s$ up to
$s=s_{c}:=\eta\left(\frac{1}{2}-f_{0}\right)\simeq\frac{\eta}{2}\,.$ (130)
@@ Appendix E Fixation probabilities in consumer-resource models
Figure S1: Theoretical and simulated fixation probabilities in consumer-resource models. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness differences $s$ by simulating mutations within consumer-resource models with (a) self-renewing resources (MacArthur consumer-resource model) and (b) externally supplied resources. In both cases, our theoretical predictions are more accurate than Kimura’s predictions, but do not fully fit with the simulation results. Error bars denote standard errors from multiple instances of demographic noise and mutants.
Figure S2: Ecological suppression of fixation probabilities controlled by species packing in consumer-resource models. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness differences $s$ by simulating mutations within MacArthur consumer-resource models with different packing fractions $S^{*}/M^{*}$. The ecological suppression of fixation probabilities is stronger when the packing fraction is lower. Error bars denote standard errors from multiple instances of demographic noise and mutants.
In the main text, we have compared our predictions on $p_{\mathrm{fix}}$ with the simulation results from generalized Lotka-Volterra model. The comparison can similarly be done for various consumer-resource models as in Figure S1. We find that our predictions are in general more accurate than Kimura’s predictions, in particular when $s<s_{c}$. On the other hand, there are some deviations after $s=s_{c}$ but not very large $s$. When $s>s_{c}$, the mutant becomes more likely to fix and the deterministic ecological dynamics affect the fixation probabilities more significantly. Compared to generalized Lotka-Volterra model, here the species interactions are mediated through resource dynamics. For self-renewing resources, there are additional complications since resources can also go extinct. We suspect that these additional complexities cause our DMFT approximations in Sec. B and C to be less accurate.
@@
Appendix F Comparison between selection and drift dynamics
@@
Both our results and the traditional results in population genetics can be qualitatively understood by comparing the relative sizes of selection and drift terms in the differential equation for $f$. As explained in the main text, in classic population genetics a mutant starts to invade at a very small frequency, where the demographic noise dominates. The mutant can either become extinct or grow to a larger frequency, both due to the demographic noise. This process continues until the surviving mutants reach a frequency $f\sim 1/N_{\mathrm{eff}}|s|$ , where the deterministic dynamics from natural selection becomes dominant instead.
Here we use similar arguments to understand the dynamics of $f$ under the presence of ecological effects. Although Eq. (43) is not exactly solvable in general, various approximations can be made such that we can compare between the sizes of selection and drift.
Suppose the mutant reaches frequency $f\gg f_{0}$ at time $t$. To compare the selection and drift effects around time $t$, it suffices to focus on small changes in frequency $\Delta f\ll f$. From Eq. (43), we can then approximate
$\frac{d(\Delta f)}{dt}\simeq f(1-f)(s-\eta f)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (132)
where $f$ is regarded as a constant in the equation. The solution is simply
$\Delta f(t+\Delta t)\simeq f(1-f)(s-\eta f)\Delta t+\sqrt{\frac{f(1-f)\Delta t}{N_{\mathrm{eff}}}}z\,$, (133)
where $z$ is a zero mean, unit variance Gaussian random variable. If drift dominates the dynamics, the above expression of $\Delta f$ remains a good approximation till
@@
To ensure that this approximation is self-consistent, the contribution to $\Delta f$ from selection must be negligible. Therefore, we require
$\left|f(1-f)(s-\eta f)\Delta t\right|\sim|s-\eta f|N_{\mathrm{eff}}f^{2}\ll f\,.$ (135)
Similarly, selection dominates the dynamics if
$|s-\eta f|N_{\mathrm{eff}}f^{2}\gg f\,.$ (136)
Therefore, we can define the selection-drift ratio
$R=|s-\eta f|N_{\mathrm{eff}}f\,$, (137)

### v2

@@ Appendix C Generalized consumer-resource models
$g_{m}\left(\vec{R}_{/pm}(t)\right)-g_{p}\left(\vec{R}_{/pm}(t)\right)\simeq\left<g_{m}\left(\vec{R}_{/pm}(0)\right)-g_{p}\left(\vec{R}_{/pm}(0)\right)\right>\,,\quad N_{p}(t)+N_{m}(t)\simeq N_{0}\,.$ (104)
@@
$\frac{df}{dt}=\,$ $f(1-f)\left(\left<g_{m}\left(\vec{R}_{/pm}(0)\right)-g_{p}\left(\vec{R}_{/pm}(0)\right)\right>-m_{m}+m_{p}\right$. (105)
@@
$=$ $\,f(1-f)\left(s_{\mathrm{inv}}-\eta f\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (107)
where the invasion fitness is
$s_{\mathrm{inv}}=$ $\,\left<g_{m}\left(\vec{R}_{/pm}(0)\right)-g_{p}\left(\vec{R}_{/pm}(0)\right)\right>-m_{m}+m_{p}$ (108)
$\,-N_{0}\sum_{\alpha\beta}(C_{m\alpha}-C_{p\alpha})\left(\left(I-P\right)Q^{-1}\right)_{\alpha\beta}E_{p\beta}$ (109)
@@ Appendix D Derivation and analysis of fixation probability
$\frac{df}{dt}=f(1-f)\left(s_{\mathrm{inv}}-\eta f\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (111)
where $s_{\mathrm{inv}}$ is the invasion fitness, $\eta$ is the strength of ecological effects depending on details of the ecological model, and $N_{\mathrm{eff}}$ is the initial effective population size of the parent and the mutant.
@@
$\frac{\partial p}{\partial t}=x(1-x)(s_{\mathrm{inv}}-\eta x)\frac{\partial p}{\partial x}+\frac{x(1-x)}{2N_{\mathrm{eff}}}\frac{\partial^{2}p}{\partial x^{2}}\,.$ (112)
@@
$0=x(1-x)(s_{\mathrm{inv}}-\eta x)\frac{\partial p^{*}}{\partial x}+\frac{x(1-x)}{2N_{\mathrm{eff}}}\frac{\partial^{2}p^{*}}{\partial x^{2}}\,$, (113)
$\Rightarrow\frac{\partial p^{*}}{\partial x}\propto\exp\left[N_{\mathrm{eff}}\left(\eta x^{2}-2s_{\mathrm{inv}}x\right)\right]\,.$ (114)
@@
$=\frac{\mathrm{Erfi}\left(\alpha\tilde{s}_{\mathrm{inv}}\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}_{\mathrm{inv}}-f_{0})\right)}{\mathrm{Erfi}\left(\alpha\tilde{s}_{\mathrm{inv}}\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}_{\mathrm{inv}}-1)\right)}\,$, (116)
@@
is the imaginary error function. We have denoted $\alpha=\sqrt{N_{\mathrm{eff}}\eta}$ as the ratio of strength between ecological effects and demographic noise, and $\tilde{s}_{\mathrm{inv}}=s_{\mathrm{inv}}/\eta$ is proportional to the “dressed invasion fitness” [14].
@@
For example, when $s_{\mathrm{inv}}=0$,
$p_{\mathrm{fix}}(s_{\mathrm{inv}}=0)=\frac{\mathrm{Erfi}\left(\alpha f_{0}\right)}{\mathrm{Erfi}\left(\alpha\right)}\,.$ (119)
@@
$p_{\mathrm{fix}}(s_{\mathrm{inv}}=0,\alpha\ll 1)\simeq\frac{\alpha f_{0}}{\alpha}=f_{0}\,.$ (120)
@@
$p_{\mathrm{fix}}(s_{\mathrm{inv}}=0,\alpha\gg 1)\sim e^{-\alpha^{2}}\,.$ (121)
@@
We see that $p_{\mathrm{fix}}(s_{\mathrm{inv}}=0)$ matches Kimura’s prediction when $\alpha\ll 1$, but exponentially suppressed by $\alpha$ otherwise.
@@
The fixation probability for neutral mutations suggests that there is a crossover between the regimes of weak and strong ecological effects, characterized by small and large $\eta$ respectively. First, for small $\eta$ with $\alpha\ll 1$, we have $|\alpha\tilde{s}_{\mathrm{inv}}|\gg 1$ for any finite $s_{\mathrm{inv}}$ since $\alpha\tilde{s}_{\mathrm{inv}}$ scales as $\eta^{-1/2}$. Therefore, we can use the exponential approximation for all the terms and $p_{\mathrm{fix}}$ becomes
$p_{\mathrm{fix}}$ $\simeq\frac{\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)-\exp\left(\alpha^{2}(\tilde{s}_{\mathrm{inv}}-f_{0})^{2}\right)}{\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)-\exp\left(\alpha^{2}(\tilde{s}_{\mathrm{inv}}-1)^{2}\right)}$ (122)
$\simeq\frac{1-\exp(-2N_{\mathrm{eff}}s_{\mathrm{inv}}f_{0})}{1-\exp(-2N_{\mathrm{eff}}s_{\mathrm{inv}})}\,$, (123)
@@
which is exactly Kimura’s formula for fixation probability with $s=s_{\mathrm{inv}}$. Combining the result for neutral mutation, we see that Kimura’s formula works equally well in ecological context as long as $\alpha\ll 1$. On the other hand, for large $\eta$ with $\alpha\gg 1$, the behavior of $p_{\mathrm{fix}}$ is divided into different regimes according to $s_{\mathrm{inv}}$. Now for $|\alpha\tilde{s}_{\mathrm{inv}}|\ll 1$, we again have
$p_{\mathrm{fix}}\simeq\frac{\mathrm{Erfi}\left(\alpha f_{0}\right)}{\mathrm{Erfi}\left(\alpha\right)}\sim e^{-\alpha^{2}}\,.$ (124)
For $|\alpha\tilde{s}_{\mathrm{inv}}|\gg 1$, we have
$p_{\mathrm{fix}}$ $\simeq\frac{\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-f_{0})\exp\left(\alpha^{2}(\tilde{s}_{\mathrm{inv}}-f_{0})^{2}\right)}{\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-1)\exp\left(\alpha^{2}(\tilde{s}_{\mathrm{inv}}-1)^{2}\right)}$ (125)
$=\frac{1-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-f_{0})\exp\left(-\alpha^{2}f_{0}(2\tilde{s}_{\mathrm{inv}}-f_{0})\right)}{1-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-1)\exp\left(-\alpha^{2}(2\tilde{s}_{\mathrm{inv}}-1)\right)}\,.$ (126)
@@
The behavior of the above expression is further divided into two regimes. When $\tilde{s}_{\mathrm{inv}}\ll 1/2$, the denominator is dominating and $p_{\mathrm{fix}}$ is still exponentially suppressed:
$p_{\mathrm{fix}}\sim\exp\left(-2N_{\mathrm{eff}}(s_{\mathrm{inv}}-\eta/2)\right)\,.$ (127)
@@
When $\tilde{s}_{\mathrm{inv}}\gg 1/2$, the denominator becomes very close to $1$ and the numerator becomes dominating instead. We then have
$p_{\mathrm{fix}}\simeq\frac{1-\exp(-2N_{\mathrm{eff}}s_{\mathrm{inv}}f_{0})}{1-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-1)\exp(-2N_{\mathrm{eff}}s_{\mathrm{inv}})}\,$, (128)
@@
which matches Kimura’s formula approximately for $\tilde{s}_{\mathrm{inv}}<1$ and exactly for $\tilde{s}_{\mathrm{inv}}>1$.
@@
In conclusion, while $p_{\mathrm{fix}}$ for negative $s_{\mathrm{inv}}$ is always exponentially suppressed regardless of ecology, the presence of a large $\eta$ now also exponentially suppresses $p_{\mathrm{fix}}$ with positive $s_{\mathrm{inv}}$ up to
$s_{\mathrm{inv}}\simeq s_{c}:=\frac{\eta}{2}\,.$ (129)
@@ Appendix E Fixation probabilities in consumer-resource models
Figure S1: Theoretical and simulated fixation probabilities in consumer-resource models. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within consumer-resource models with (a) self-renewing resources (MacArthur consumer-resource model) and (b) externally supplied resources. In both cases, our theoretical predictions are more accurate than Kimura’s predictions, but do not fully fit with the simulation results. Error bars denote standard errors from multiple instances of demographic noise and mutants.
Figure S2: Ecological suppression of fixation probabilities controlled by species packing in consumer-resource models. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within MacArthur consumer-resource models with different packing fractions $S^{*}/M^{*}$. The ecological suppression of fixation probabilities is stronger when the packing fraction is lower. Error bars denote standard errors from multiple instances of demographic noise and mutants.
In the main text, we have compared our predictions on $p_{\mathrm{fix}}$ with the simulation results from generalized Lotka-Volterra model. The comparison can similarly be done for various consumer-resource models as in Figure S1. We find that our predictions are in general more accurate than Kimura’s predictions. On the other hand, there are some deviations for $0<s_{\mathrm{inv}}<\eta$, where the parent and the mutant can transiently coexist, hence the deterministic ecological dynamics affect the fixation probabilities more significantly. Compared to generalized Lotka-Volterra model, here the species interactions are mediated through resource dynamics. For self-renewing resources, there are additional complications since resources can also go extinct. We suspect that these additional complexities cause our DMFT approximations in Sec. B and C to be less accurate.
@@
Appendix F Effects of parent-mutant correlation on the mutation dynamics
Figure S3: Dynamics of parent and mutant abundances during fixation of a neutral mutation at various parent-mutant correlation. Fixation of a neutral mutation requires deviation from the deterministic ecological steady state. At high $\rho=0.9$ (a), the total parent-mutant abundance remains constant on average throughout the full dynamics. At low $\rho=0.5$ (b), the total parent-mutant abundance drops significantly towards the final mutant abundance.
Figure S4: Theoretical and simulated fixation probabilities in generalized Lotka-Volterra models at various parent-mutant correlation. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within generalized Lotka-Volterra models. Our theory (solid curves) matches the simulations (dots) for high $\rho$, but deviates from the simulations for low $\rho$ (a). The relative error is more significant for lower $s_{\mathrm{inv}}$ (b). Error bars denote standard errors from multiple instances of demographic noise and mutants.
Throughout the paper, we have assumed high parent-mutant correlation $\rho\approx 1$. In fact, this is one of the crucial assumptions that allows us to simplify the mutation dynamics into the form in Eq. (42). Here we evaluate in more detail how parent-mutant correlation affects the mutation dynamics.
Qualitatively, when the parent and the mutant are highly similar in their ecological interactions with the community, the total parent-mutant population behaves effectively as a single species from the community perspective. Therefore, the community as a whole stays near the ecological steady state. This is true even when the parent and the mutant deviate much away from their steady states due to demographic noise; such deviation is required for a mutant with low $s_{\mathrm{inv}}$ to fix and replace the parent.
The fixation dynamics simplifies in the vicinity of the community steady state in the following ways. First, the total parent-mutant abundance $N_{p}+N_{m}$ always stays constant on average, with fluctuations coming from demographic noise only. In contrast, when $\rho$ is not high, $N_{p}+N_{m}$ is not stationary on average when a mutant with low $s$ fixes, as shown in Fig. S3. Second, the community response throughout the dynamics can be approximated by the one at the steady state. Correspondingly, the susceptibility $\nu(t,t^{\prime})$ can be replaced by $\nu\delta(t-t^{\prime})$ as predicted using the cavity method for the steady state. As shown in Fig. S4, this approximation for the DMFT indeed becomes less accurate when $\rho$ is not high, leading to errors in fixation probabilities. We suspect that a more careful treatment of the DMFT can predict mutation outcomes without high $\rho$.
Appendix G Comparison between selection and drift dynamics
In this appendix, we obtain a qualitative understanding of both the results in classic population genetics and our new results, by comparing the relative sizes of selection and drift terms in the differential equation for $f$. This analysis helps us understand how parent-mutant coexistence arises in the presence of ecological effects.
G.1 Without ecological effects
@@
For pedagogical purposes, it is helpful to start by re-deriving some classic results from population genetics in the absence of ecology. As explained in the main text, a mutant starts to invade at a very small frequency where genetic drift dominates. The mutant can either become extinct or grow to a larger frequency due to drift. This process continues until the surviving mutants reach a frequency $f\sim 1/N_{\mathrm{eff}}|s|$ , where selection becomes dominant instead. This critical frequency is the drift threshold as shown in Fig. 4a in the main text.
We now obtain a heuristic analytical estimate of drift threshold, following the approach in [80, 81]. Recall that the dynamics of $f$ in classic population genetics is given by
$\frac{df}{dt}=f(1-f)s+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,.$ (131)
The exact solution to this differential equation is in general complicated as both the growth rate and diffusion rate depends on $f$. On the other hand, analogous to the Euler method for numerically solving differential equations, we can approximate the dynamics of $f$ by splitting it into infinitesimal time intervals with width $\Delta t$. Within a time interval around some time $t$, the changes in both the growth rate and the diffusion rate can be neglected. Therefore, the change in frequency $\Delta f$ within the time interval can be easily solved and is given by
$\Delta f:=f(t+\Delta t)-f(t)\simeq f(t)(1-f(t))s\Delta t+\sqrt{\frac{f(t)(1-f(t))\Delta t}{N_{\mathrm{eff}}}}z\,$, (132)
where $z$ is a zero mean, unit variance Gaussian random variable. Below, we can simply use $f$ to denote $f(t)$.
Let us focus on small frequencies $f\ll 1$ where drift dominates over selection. In this case, $\Delta f$ can be either positive or negative due to randomness, and its magnitude can be estimated by the variance
$|\Delta f|\sim\sqrt{\frac{f(1-f)\Delta t}{N_{\mathrm{eff}}}}\,.$ (133)
The above approximation of constant growth and diffusion rates remains valid until $\Delta f$ is comparable to $f$, that is
@@
Within this time interval, the contribution to $\Delta f$ from selection is
$|\Delta f|_{\mathrm{sel}}\sim f(1-f)|s|\Delta t\sim|s|N_{\mathrm{eff}}f^{2}\,.$ (135)
@@
Note that the absolute value is needed since $\Delta f$ due to selection can also be either positive or negative depending on the sign of $s$. To ensure that the approximation is self-consistent, $\Delta f$ due to selection must be negligible, i.e., $|\Delta f|_{\mathrm{sel}}\ll|\Delta f|$. Therefore, we require
$|s|N_{\mathrm{eff}}f^{2}\ll f\Rightarrow f\ll\frac{1}{N_{\mathrm{eff}}|s|}\,.$ (136)
Indeed, we see that drift dominates the dynamics when $f$ is below drift threshold.
G.2 With ecological effects
We can repeat the above analysis under the presence of ecological effects. As explained in the main text, when $0\lesssim s_{\mathrm{inv}}\lesssim\eta$, there is an additional deterministic fixed point $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$. We then expect a new region around both sides of $f^{*}$ where drift dominates the dynamics, apart from the one near $f=0$. This is because at every fixed point, by definition, the deterministic contribution to the dynamics vanishes.
Suppose the mutant reaches frequency $f\gg f_{0}$ at time $t$. Again, we focus on an infinitesimal time interval with width $\Delta t$. From Eq. (42), we can approximate $\Delta f$ within this time interval as
$\Delta f\simeq f(1-f)(s_{\mathrm{inv}}-\eta f)\Delta t+\sqrt{\frac{f(1-f)\Delta t}{N_{\mathrm{eff}}}}z\,$, (137)
where $z$ is a zero mean, unit variance Gaussian random variable. Note that the variance is the same as in the case without ecological effects. Therefore, if drift dominates the dynamics, the above expression of $\Delta f$ remains a good approximation until
$\Delta t\sim\frac{N_{\mathrm{eff}}f}{1-f}\,$, (138)
which is the same as Eq. (134). To ensure that this approximation is self-consistent, the contribution to $\Delta f$ from selection must be negligible. Therefore, we require
$\left|f(1-f)(s_{\mathrm{inv}}-\eta f)\Delta t\right|\sim|s_{\mathrm{inv}}-\eta f|N_{\mathrm{eff}}f^{2}\ll f\,.$ (139)
Note again that the absolute value is needed since frequency can either increase or decrease, i.e., $\Delta f$ due to selection can be either positive or negative, particularly when $f$ is around the new fixed point $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$, which is between $f=0$ and $f=1$. Based on the above condition, we can define the selection-drift ratio
$R=|s_{\mathrm{inv}}-\eta f|N_{\mathrm{eff}}f\,$, (140)